If a proton and an electron are released when they are 5.50×10−10 m apart (typical atomic distances), find the initial acceleration of each of them.
Answer #1

We know the initial force of atrction F = kq1q2 / r^2

where k is the Coulomb constant, and q1 and q2 are actually the same charge, that of the electron and proton e = 1.6x10^-19 C and r is the distance separating the particles.


We also know that F = ma

so we can equate the two terms to get

Accelerattion of electron

meae = ke^2 / r^2
or
ae = ke^2 / mer^2 = [9 x 109 x (1.6 x 10-19)2] / [9.1 x 10-31 x (5.5 x 10-10)2]

ae = 8.4 x 1020 m/s2

Accelerattion of proton

mpap = ke^2 / r^2
or
ap = ke^2 / mpr^2 = [9 x 109 x (1.6 x 10-19)2] / [1.67 x 10-27 x (5.5 x 10-10)2]

ap = 4.6 x 1017 m/s2